A loan of ​$2000 at 12​% is compounded monthly for two years. Find the future value and compound interest. Use the​ $1.00 future value table or the future value and compound interest formula. Click here to view page 1 of the future value table.LOADING... Click here to view page 2 of the future value table.LOADING... Click here to view page 3 of the future value table.LOADING... Click here to view page 4 of the future value table.LOADING...

The future value of the loan is ​$ . ​(Round to the nearest cent as​ needed.)

The compound interest is ​$ .​ (Round to the nearest cent as​ needed.)
Answer #1

Solution:

Monthly interest rate = 12%/12 = 1%

Nos of monthly periods = 2*12 = 24

Future value of loan = $2,000 * (1+0.01) ^ 24 = $2,000 * 1.26973 = $2,539.46

Compound interest = $2,539.46 - $2,000 = $539.46